OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + x - 30

OpenStudy (anonymous):

choices

OpenStudy (anonymous):

A. x - 6 B. x + 3 C. x - 3 D. x + 6

OpenStudy (gorv):

x2+6x-5x-30=0 x(x+6)-5(x+6)=0 (x+6)(x-5)=0

OpenStudy (gorv):

@youngstar10